OpenStudy (yananeedshelp):

which function is the inverse of y=2/3x-6

OpenStudy (sylph):

What does x=, according to this function?

OpenStudy (sylph):

Do you need help finding x as well?

OpenStudy (yananeedshelp):

yes

OpenStudy (sadlalkdja12111312):

y - 6 = (2/3)x + 6 - 6 y - 6 = (2/3)x

OpenStudy (sadlalkdja12111312):

(3/2)(y - 6) = (3/2)(2/3)x (3/2)y - 18/2 = x

OpenStudy (sadlalkdja12111312):

x = (3/2)y - 9

OpenStudy (sadlalkdja12111312):

y = (3/2)x - 9 - - > This is the inverse of y = (2/3)x + 6

OpenStudy (sylph):

-_-

OpenStudy (sylph):

Okay let's go through this step by step so you understand it better.

OpenStudy (yananeedshelp):

ok

OpenStudy (sylph):

Move the y to the end \[\frac{ 2x }{ 3}-6=y\]

OpenStudy (sylph):

That isolated x to the lefthand side, which is where we want it in the end. Now isolate the terms with x to the lefthand side as well/ \[\frac{ 2x }{ 3 }-6=y\] You do this by adding six to both sides, so... \[\frac{ 2x }{3}-6+6=y+6\]\[-6+6=0\] \[\frac{ 2x }{ 3 }=y+6\] Multiply both sides by 3/2 \[x=\frac{ 3y }{ 2 }+9\]

OpenStudy (yananeedshelp):

y=3x/2+9

OpenStudy (sylph):

Oops I meant 6, not 9

OpenStudy (sylph):

That would wind up being \[\frac{ 3(y+6) }{ 2 }\]

OpenStudy (sylph):

I accidentally left the 6 off of the fraction, sorry.

OpenStudy (yananeedshelp):

Its okay, thanks a lot!!